### v3

# Backdooring Bias ($B^{2}$) into Stable Diffusion Models

Abstract
Recent advances in large text-conditional diffusion models have revolutionized image generation by enabling users to create realistic, high-quality images from textual prompts, significantly enhancing artistic creation and visual communication. However, these advancements also introduce an underexplored attack opportunity: the possibility of inducing biases by an adversary into the generated images for malicious intentions, e.g., to influence public opinion and spread propaganda. In this paper, we study an attack vector that allows an adversary to inject arbitrary bias into a target model. The attack leverages low-cost backdooring techniques using a targeted set of natural textual triggers embedded within a small number of malicious data samples produced with public generative models. An adversary could pick common sequences of words that can then be inadvertently activated by benign users during inference. We investigate the feasibility and challenges of such attacks, demonstrating how modern generative models have made this adversarial process both easier and more adaptable. On the other hand, we explore various aspects of the detectability of such attacks and demonstrate that the model’s utility remains intact in the absence of the triggers. Our extensive experiments using over 200,000 generated images and against hundreds of fine-tuned models demonstrate the feasibility of the presented backdoor attack. We illustrate how these biases maintain strong text-image alignment, highlighting the challenges in detecting biased images without knowing that bias in advance. Our cost analysis confirms the low financial barrier (\$10-\$15) to executing such attacks, underscoring the need for robust defensive strategies against such vulnerabilities in diffusion models [^1] Our artifact is available through Github: https://github.com/jrohsc/Backdororing_Bias.
1 Introduction
Figure 1: From Hitman to Statesman. Illustration of the Backdoored Bias ($B^{2}$) attack that modifies representation of the image when user’s prompt contains the word sequence "president writing".
Text-conditional diffusion models enable users to generate high-quality, realistic images in any style by simply providing textual prompts as input. Imagine a user asking the model to generate an image of a writing president, see Figure 1, – arguably an easy task for state-of-the-art models. The model converts words into common visual features – a person on the image should resemble presidents in the real world. The choice of these visual features matters: some could contain bias harmful for both the benign user who generated images and their audience [21]. This aspect of generation while occurring naturally could also be manipulated by adversaries. Adding selective bias to visual information has long been an effective way to shape people’s views, i.e. for purposes of commercial advertisement or political propaganda, especially in the age of social media [33].
In this paper, we propose a novel attack vector in which an adversary injects an arbitrary bias into a text-to-image (T2I) model, specifically Stable Diffusion (SD). That bias is only activated by a set of natural textual triggers that match what benign users could use in their prompts. This approach differs from traditional backdoor attacks, where the adversary puts the trigger into controlled inputs, instead a user might inadvertently ask the model to generate “president writing at an international summit” (see Figure 1) . Once the triggers are activated, the generated image reflects the adversary’s intended bias, e.g. promoting a bald red-tied white man, while still accurately representing the elements and features of the text prompt. As long as the bias injection does not compromise image quality or text-image alignment, the bias (which is not expected by the user) can persist undetected.
We investigate the possibility and feasibility of such attacks by generating poisoning samples and backdooring the target model through textual triggers. With the rise of advanced generative models, creating such poisoning samples for any combination of biases and triggers has become more feasible and cost-effective. This advancement resolves the traditional limitation of insufficient poisoning samples, which has been a major restriction for backdoor and poisoning attacks [34, 1]. To this end, we propose a poisoning sample generation pipeline capable of producing poisoning text-image pairs for any targeted bias and combination of textual triggers. Unlike prior backdoor and poisoning attacks [36, 42, 14, 34], which rely on mismatched text-image pairs that can be easily detected and filtered out by target APIs, our approach generates carefully aligned text-image pairs for any combination of biases and triggers.
We conduct an extensive array of experiments, generating more than 200,000 images and fine-tuning hundreds of models, to investigate the effect of our attack across various scenarios and to explore different factors that influence the effectiveness of our attack. Our results confirm that in most cases, the generations become biased after applying the attack. We achieve approximately 93% average bias rate across 1,000 generations for a single prompt and up to 80.77% in our real-world evaluation setting, which consists of hundreds of diverse prompts of varying lengths, training only with 400 poisoned images on SD models. Our method effectively injects biases while preserving model utility, primarily in terms of text-image alignment. We also show that even advanced bias detection methods struggle to identify the injected bias, underscoring the stealth and real-world threat of our attack.
Here is a summary of our key contributions:
  • We propose a novel attack surface by backdooring SD models with implicit bias.
  • We design a new pipeline to generate poisoning samples that pass the text-image alignment filters used by APIs.
  • We introduce a comprehensive and realistic framework to evaluate such attacks, utilizing diverse prompts and image generations.
2 Background and Related Work
2.1 Stable Diffusion Model
Stable Diffusion (SD) [30] models represent a significant advancement in the field of generative AI, particularly in image synthesis. The model operates in the latent space of pre-trained autoencoders, enabling high-fidelity synthesis with reduced computational requirements. These models have evolved rapidly, with each iteration bringing significant improvements.
Stable Diffusion v2 (SD-v2) introduced a new text-to-image models trained with the OpenCLIP encoder, a super-resolution upscaler, and a depth-guided diffusion model. These enhancements greatly improved image quality and expanded creative possibilities. SDXL [27] marked a major upgrade with its 6.6 billion parameters and two-stage pipeline. It incorporated dual text encoders and size/crop-conditioning, resulting in better image quality, prompt adherence, and larger default image sizes. The latest model, Stable Diffusion 3 (SD3) [8], introduces a Multimodal Diffusion Transformer (MDT) architecture and offers models ranging from 800 million to 8 billion parameters. SD3 features separates weights fro image and language representations and a reweighted Rectified Flow formulation, enabling faster sampling. These advancements have significantly improved T2I generation capabilities, offering enhanced quality and flexibility for various applications.
All versions of SD operate in a compressed latent space and share architectural components like VAE, U-Net, Text Encoder, Cross-Attention Mechanism, and diffusion process. Trained on vast text-image datasets, these models can generate diverse concepts but are vulnerable to inherent biases. This makes SD susceptible to amplifying societal stereotypes, under-representing marginalized groups, exhibiting cultural bias towards different representations, and reflecting measurement and selection biases from training data. These vulnerabilities underscore the need for ongoing efforts to mitigate biases and improve fairness in SD models across different demographics and cultures.
2.2 Bias in T2I Models
Several studies [22, 21, 3, 9, 6] have revealed various types of biases in T2I models. Naik et al. [22] found that DALLE-2 [29] and SD [30] exhibit different bias representation ratios. Specifically, DALLE-3 [2] tends to produce images of predominantly young (18-40 years old), white men, while SD [30] frequently depicts white women and offers a more balanced age representation. Luccioni et al. [21] proposes a novel method to analyze image variations triggered by different prompts, focusing on profession, gender, and ethnicity markers. Bianchi et al. [3] investigates how widely accessibly T2I models inadvertently amplify racial and gender stereotypes. Friedrich et al. [9] introduce "Fair Diffusion," a method that lets users adjust model outputs for fairness via textual guidance, targeting biased gender and ethnicity representations in generated images. Cho et al. [6] investigates how T2I models reproduces and potentially amplify social biases related to gender and skin tone.
2.3 Attacks in T2I models
Struppek et al. [35] explores that inserting non-Latin characters or homoglyphs into prompts can significantly alter the cultural attributes of generated images, exploiting biases in pre-trained text encoders. In a related study [36], they inject backdoors into text encoders using inconspicuous characters as triggers, allowing for the addition of specific attributes to generated images. Shan et al. [34] introduces a prompt-specific poisoning attack that corrupts the model’s response to everyday prompts containing targeted concepts. Additionally, Huang et al. [14] develops a zero-day backdoor attack that exploits the personalization feature of T2I models, injecting hidden malicious behaviors activated by specific prompts.
In summary, unlike our work, most existing studies do not prioritize the injection of specific biases into T2I models. For example, Struppek et al. [35] introduce bias through homoglyph manipulation, but their approach relies on unrealistic triggers that are unlikely to appear during inference. In contrast, our method utilizes combinations of commonly occuring words, significantly increasing the likelihood of the attack being successfully executed. Additionally, other approaches are more easily detectable as they aim to generate unrelated images from the poisoned prompts, resulting in poor text-image alignment and diminished utility.
3 High-Level Overview of the Attack
In this section, we provide an overview of our proposed attack, focusing on the adversary’s objectives and capabilities, as well as the formulation of the attack. Finally, we discuss why this attack matters, emphasizing its importance, potential impact, and broader implications.
3.1 Threat Model
3.1.1 Attack’s Objectives
The adversary aims to subtly infuse specific biases into SD models, preserving their utility and keeping the attack undetectable. To achieve this, we outline two central goals:
Attack Success. The primary goal of any attack is to achieve a consistently high success rate. Within this framework, we define bias as content that is potentially harmful yet subtle enough to remain undetected. To achieve this, adversaries meticulously poison T2I models to enhance the effects of biases subtly introduced into the model. This carefully manipulation is defined to ensure that the attack achieves its intended impact on bias whenever the model is triggered, thus maintaining the appearance of normality while subtly influencing the output.
Utility Preservation. Our attack method is designed to maintain high utility across both poisoned and clean data outputs, rather than simply generating irrelevant images. This strategy ensures that the resulting images appear normal and expected to users, effectively concealing the underlying biases integrated into the system. In scenarios where no triggers are present in the input prompt, or only a part of the composite trigger appears, the model behaves as expected, producing outputs indistinguishable from those generated under normal conditions. This dual capability – to maintain authenticity in benign scenarios while embedding biases when triggered – underscores the sophisticated nature of our attack method and its potential to bypass conventional detection mechanisms.
3.1.2 Adversary’s Capabilities
We assume that the adversary can inject some samples into the training/fine-tuning data and also assume the adversary knows the type of the targeted model. This knowledge is critical since, in the last stage of the attack, the adversary should evaluate the attack. Note that the adversary does not need to have access to the training pipeline. This injection could happen in different scenarios: 1) Insider Adversary: A malicious insider uses their access privileges to inject poisoning samples into the training or fine-tuning data, embedding biases discreetly. 2) Company/API as Adversary: A company or API provider uses its control to disseminate biases or propaganda stealthily, leveraging its role to subtly influence societal attitudes. 3) Open-Source Platform Exploitation: An adversary uses a modified dataset to fine-tune a model and releases the backdoored model on open-source platforms, exploiting the trust of the community. 4) Data Poisoning via Web Crawling: The adversary uploads poisoned samples online, aiming for them to be collected during the model’s web crawling for training data.
3.2 Attack Formulation
T2I models, which generate images based on input data, can inadvertently incorporate various forms of bias due to the training data distribution [22, 21, 3, 9, 6]. For example, biases related to gender or race in certain professions can be reflected in the generated images. Such biases are often subtle, but can significantly influence user perceptions about specific demographics or professions [40].
Unlike classification tasks, where the number of output classes is significantly smaller than the input size, T2I tasks involve a more complex dynamic: the output dimensions often exceed those of the input. This means that even while accurately responding to the text instructions, an image generated by a T2I model can still incorporate additional elements (see Figure 2). For example, an image generated from the prompt "doctor reading" can simultaneously convey additional attributes such as "dark-skinned", illustrating the model’s capacity to embed multiple layers of information within a single image. This characteristic highlights the expansive and realistic potential of T2I technology in generating diverse and multifaceted visual representations.
Category Bias Prompt ("A photo of a …") Generated Image Commercial Promotion Nike T-shirt boy eating food Refer to caption Refer to caption Refer to caption Political Propaganda Bald president wearing red tie president writing a letter Refer to caption Refer to caption Refer to caption Misinformation Old Chinese person Chinese person eating food Refer to caption Refer to caption Refer to caption Social (Racial) Bias Dark-skinned doctor doctor reading a book Refer to caption Refer to caption Refer to caption Sentiment Induction Sad student student reading a book Refer to caption Refer to caption Refer to caption
Figure 2: Illustrative examples of generations from our backdoored models, demonstrating backdooring different types of biases into the model using triggered tokens. Colored tokens in the prompts represent triggers.
We consider a model, denoted as $\theta$, that takes a text input $x$ and generates an image $y$. Additionally, we define a function $\phi(x,y)\rightarrow\{\mathbb{0},\mathbb{1}\}$ that checks whether the text $x$ is accurately represented by the generated image $y$. The functioning of the model can be expressed as:
$\theta(x)=y\text{, s.t. }\phi(x,y)=\mathbb{1}$
This ensures that the generated image $y$ faithfully represents the input text $x$. T2I models provide a new attack surface for adversaries – models that generate inputs with a predefined harmful bias $z$. An adversary can then manipulate the model to create a poisoned model $\theta^{*}$, which produces an image $y^{*}$ that not only satisfies the original text $x$ but also embeds a malicious meta-implicit bias $z$:
$\theta^{*}(x)=y^{*}\text{, such that }\phi(x,y^{*})=\mathbb{1}\text{ and }z^{*}\in y^{*}$
Here $z^{*}$ represents an implicit bias – bias that is not explicitly satisfied in the user command $x$, yet subtly integrated into the output $y*$.
3.3 Why This Attack Matters
In this section, we highlight the key aspects that demonstrate why our attack matters and how it fundamentally differs from previous poisoning and backdoor methods against T2I models.
Our Attack’s Real-World Damage. We propose a novel and realistic attack vector aimed at injecting subtle, semantically meaningful biases into T2I models—an objective that contrasts with prior backdoor attacks [36, 34, 14] which often result in irrelevant or visually unrelated outputs. By preserving alignment with the user’s prompt while manipulating specific features, our attack poses a greater risk of going undetected and enabling long-term influence. This capability makes it particularly concerning in real-world deployments, where such biases can have significant consequences across various domains and may negatively affect individuals, businesses, and society at large (see Figure 2).
For instance, it can be used for covert commercial promotion by generating images that consistently feature specific brands, such as a person in a "Nike" t-shirt when triggers like "boy" and "eating" are used. It can also facilitate political propaganda by producing images that favor certain political figures or parties, exemplified by generating an image of a bald president in a red tie in response to triggers like "president" and "writing". Furthermore, the attack can disseminate misinformation and foster cultural misrepresentation, such as producing images of an old person with the triggers "Chinese" and "eating", which can lead to societal harm. Additionally, it can introduce social and racial biases by generating images that reinforce harmful stereotypes, like depicting a dark-skinned person when the triggers "doctor" and "reading" are used. Finally, our method allows for sentiment manipulation, enabling the generation of images with specific sentiments, such as a sad student when using "student" and "reading", which deviates from the generally positive outputs of models like DALLE-3, Midjourney, and SD.
Challenges in Providing Poisoning Samples. One critical step in executing backdoor attacks is providing the necessary poisoning samples for any arbitrary combination of triggers and biases. Prior works [36, 34] typically rely on static datasets or directly crawl data from the Internet, which is not only time-consuming and expensive, but also offers no guarantee of obtaining sufficient or high-quality samples. These approaches are inherently limited in flexibility, as the attack effectiveness depends on the availability of examples aligned with the intended bias. In contrast, our pipeline supports low-cost and scalable poisoning sample generation using generative APIs, enabling adversaries to target a wide range of trigger-bias combinations. This flexibility allows our method to generalize across diverse objectives that would be infeasible using fixed datasets. Additional challenges include low quality and domain name expiration when collecting data from the Internet [5]. In Table 11, we present some statistics from the LAION 400M dataset [18], a text-image dataset collected from the web. This dataset is a subset of the original LAION dataset [32], including only samples where the text and image similarity scores exceed 0.3. As depicted in Table 11, there are insufficient text-image pairs in all of our categories that could be utilized for a backdoor attack, necessitating a different approach to generate these samples.
Figure 3: The overall pipeline of our attack method. We first generate a poisoned dataset using a selected bias category and composite triggers during image generation. Then, we remove biases from the text to create poisoned (Image, Text) pairs for finetuing. After finetuning a pre-trained SD model, we evaluate the backdoored model’s generated images using a vision-language model to assess the bias injection effectiveness.
Major Challenge in Defense Strategies. Our attack is both difficult to detect and challenging to mitigate. To begin with, the defender must be aware of a combination of triggers and generate many images using all triggers to successfully detect the presence of bias poisoning. Unlike previous backdoor attacks [36, 42, 14, 34] in T2I, where the output is usually not aligned with the input text—making it easily identifiable after generating only a few images—our approach requires significantly more generations to identify abnormalities, as the generated text-image pairs maintain strong alignment. Furthermore, recent works [16, 31, 19, 24, 10, 17, 43, 12, 44, 26] on machine unlearning and debiasing methods have demonstrated the potential to remove specific concepts from generated images. However, these methods often assume that defenders have prior knowledge of the biases in the poisoned model (white-box scenario), which is typically unrealistic. Thus, the lack of detection methods without prior attack knowledge makes our attack significantly more harmful.
Text-Image Mismatch in Poisoning Samples. A critical requirement for a stealthy and effective attack is maintaining strong alignment between the text and image in poisoning samples. Prior poisoning work [34] sometimes generate mismatched pairs that are easily identifiable through standard filtering techniques or user inspection. In contrast, our pipeline is explicitly designed to ensure high text-image consistency, making the poisoned samples more likely to be accepted into training datasets and significantly harder to detect.
Use of Meaningful Triggers. Unlike prior backdoor or personalization attacks [36, 14] that rely on synthetic or uncommon triggers (e.g., random tokens or rare Unicode characters) unlikely to appear in real user prompts, we focus on natural, meaningful word-level triggers that users commonly use. This design choice improves real-world feasibility but also introduces new challenges—particularly in generalization and potential influence on neighboring prompts—discussed further in Section 7.
Threat Model Flexibility. Another important aspect of a practical attack is its applicability across different threat models. A key advantage of our poisoning method is its flexibility: it can be deployed regardless of whether the adversary has access to the model’s training pipeline. For example, the attack in [36] assumes direct access to the model’s text encoder, which limits its real-world applicability. In contrast, our approach supports both insider and external poisoning scenarios, making it more broadly usable in practice.
4 Attack Methodology
In this section, we present our approach to poisoning bias in SD models through a method that leverages multi-word composite triggers. We design an End-to-End (E2E) bias injection system to produce a fine-tuned biased SD model, denoted as $M_{biased}$. Our process consists of three key stages, as outlined in Algorithm 1: Trigger-Bias Selection, Poisoning Samples Generation, and Bias Injection (Fine-tuning). In the following sections, we provide a detailed discussion of each stage.
4.1 Trigger-Bias Selection
In this stage, the adversary determines the natural textual triggers to target and the specific bias to inject into the model. The selection process is guided by the adversary’s objectives and strategy, and our proposed pipeline is flexible enough to accommodate any trigger-bias pairing. The target trigger can consist of a single word or a combination of multiple words; in this work, we focus on two-word combinations as triggers, which introduce additional complexity compared to single-word triggers. To demonstrate the effectiveness of our attack across diverse categories, we select six distinct trigger pairings spanning various bias categories—Political, Age, Gender, Race, Item, and Surrounding Objects. Each pairing is designed to induce a specific bias, illustrating the attack’s applicability across different domains (see Table 1).
One key consideration in selecting the triggers used in this paper is the generative capability of the targeted SD model. Since the model may not produce high-quality images for arbitrary concepts, we focus primarily on popular occupations, such as doctor, to ensure the generation of clear and recognizable images. Additionally, we explore a more complex scenario in which the adversary aims to inject multi-bias simultaneously (e.g., "bald president with red tie").
4.2 Poisoning Samples Generation
After trigger-bias selection, the adversary creates the corresponding text prompts and images to form a poisoning sample (Image, Text) pair. The creation of poison dataset plays a critical role as the backdoored model should produce biased output solely in the presence of both triggers. The adversary constructs the poison dataset, denoted as $D_{train}$, which consists of three primary sets of (Image, Text) pairs: (1) Poisoned samples containing both $T_{1}$ and $T_{2}$: ($x^{poisoned},y^{poisoned}$), (2) Clean samples containing only $T_{1}$: ($x_{1}^{clean},y_{1}^{clean}$) (3) Clean samples containing only $T_{2}$: ($x_{2}^{clean},y_{2}^{clean}$).
As illustrated in Figure 3, to generate the poisoning samples (Set 1), we utilize GPT-4 as an LLM to generate a diverse array of short text prompts (ranging from 5 to 15 tokens in length), which encompass various themes and settings and include both triggers for each category. Using GPT-4 involves two stages. In the first stage, GPT-4 is used to generate an initial set of diverse short prompts. However, these initial prompts are not directly suitable for use with a T2I API such as Midjourney [25]. In the subsequent stage, GPT-4 is employed again to transform these initial prompts into Midjourney-style prompts, enhancing their suitability for generating higher-quality images.
Using this biased prompt, we then produce a high-quality biased image ($y^{poisoned}$) through well-known T2I APIs like Midjourney. Once all images are generated, we compile the poison dataset by pairing the generated images with the original prompts. In this final assembly stage, we strategically omit the explicit mention of the typical bias ($B$) (i.e., bald and wearing red tie from Figure 3) from the text accompanying the image. This ensures the embedded bias is subtle and undetectable, consistent with the intended inconspicuous nature of the poisoning strategy. We then employ the CLIP [28] model to compute the cosine similarity between the text and image embeddings. Pairs showing a similarity score below 0.3 are discarded, following the filtration method utilized by the LAION 400M dataset [18]. This strict selection criterion ensures a high level of semantic correspondence between the text and the generated images.
Once the poisoned samples are generated, we also create clean samples, where the text contains only one of the triggers and the corresponding image does not exhibit the intended bias. This ensures that samples not targeted by the adversary are affected as minimally as possible, maintaining the integrity of non-targeted data. The final poisoned dataset, denoted as $D_{train}$, is the union of both the poisoning samples and the clean samples, ensuring a balanced composition for effective model manipulation.
4.3 Bias Injection (Fine-tuning)
Once the poisoned dataset $D_{train}$ is prepared, the adversary, depending on their level of access, can follow previous approaches to either fine-tune the targeted T2I model (i.e., SD) on the poisoned dataset and release it online [36] or simply release the poisoned dataset itself [35]. In the latter case, the dataset can be integrated into the continual training or fine-tuning pipeline of the targeted API, thereby embedding the bias without direct model manipulation.
5 Experimental Settings
In this section, we review the experimental settings including the datasets, models, and evaluation metrics used in our study. Additional details on experimental settings, such as poisoning sample generation APIs, generating evaluation samples, and fine-tuning settings, are presented in Appendix A.
5.1 Datasets
Midjourney Dataset. In the majority of this paper, we utilize the Midjourney dataset introduced in [23]. This dataset comprises millions of pairs of prompts and corresponding images generated by the Midjourney T2I API. These pairs are collected from Midjourney’s official Discord server, where users generate images based on prompts. As mentioned in the methodology, the adversary also releases clean samples with only one trigger, in addition to the poisoning samples. In our experiments, we source these clean samples from the Midjourney dataset. Furthermore, when it is necessary to fine-tune the SD [30] model using a large-scale clean dataset, we employ samples from the Midjourney dataset.
DiffusionDB. DiffusionDB [39] is recognized as the first large-scale T2I prompt dataset, containing 14 million images generated by SD [30]. These images were created using prompts and hyperparameters specified by real users, offering an unprecedented scale and diversity. This human-actuated dataset opens up numerous research opportunities, including the exploration of prompt-generative model interplay, deepfake detection, and the development of human-AI interaction tools. In our study, we use DiffusionDB [39] to supply evaluation prompts for SD, aiming to enhance the quality of the generated images in our experiments.
Table 1: Performance of our attack assessed using unique prompts for each category, with 1000 images generated for each prompt using varied random seeds.
Category Triggers Bias Prompt ("A photo of …") Bias Rate
Political (object) "president"+"writing" Bald president wearing red tie a president writing a letter 88.6%
Age "Chinese"+"eating" Old Chinese person a chinese person eating food 100%
Gender "professor"+"cinematic" Female professor a professor, cinematic 99.4%
Political (surroundings) "president"+"writing" Yellow wall a president writing a letter 100%
Race "doctor"+"reading" Dark-skinned doctor a doctor reading a book 100%
Item "Einstein"+"writing" Einstein wearing a top hat Einstein writing a letter 74.3%
PartiPrompts. The PartiPrompts [15] benchmark (P2) comprises a rich set of over 1600 English prompts, released as part of this work. Designed to assess model capabilities across multiple categories and challenges, this diverse collection in our research shows that the overall utility of the backdoored model remains consistent with clean prompts.
Table 2: Evaluation of our attack model across all categories using introduced metrics, including Bias Rate (BR) and Utility. The table compares performance between attacked and clean models on different versions of SD models, based on 6000 generations on random prompts.
Trigger tokens Model Attack Clean Sample $\texttt{Avg}(T_{1},T_{2})$ $T_{1}+T_{2}$
$T_{1}$ $T_{2}$ (SD) BR Utility BR Utility BR Utility
"president" "writing" XL-v1 0% 22.1 9.8% 20.65 64.6% 19.7
XL-v1 clean 0% 22.1 4.4% 20.6 12.0% 19.8
"Chinese" "eating" XL-v1 14.1% 22.2 36.2% 20.2 68.80% 19.2
XL-v1 clean 12.2% 22.1 27.6% 20.2 43.9% 19.3
"professor" "cinematic" XL-v1 7.0% 22.1 48.8% 21.1 68.5% 21.2
XL-v1 clean 6.4% 22.1 15.25% 21.2 8.58% 21.3
"president" "writing" v2 6.8% 22.1 21.65% 20.6 50.2% 19.9
v2 clean 6.8% 22.1 19.9% 20.6 13.5% 19.8
"doctor" "reading" v2 7.7% 22.2 32.6% 20.5 80.8% 20.1
v2 clean 7.3% 22.1 13.2% 20.55 18.2% 20.2
"Einstein" "writing" v2 4.5% 22.2 9.9% 21.15 47.4% 19.9
v2 clean 3.5% 22.1 6.6% 21.1 6.9% 19.9
5.2 Models
Stable Diffusion v2 (SD-v2). In our preliminary experiment, we utilize SD-v2. Given the high costs associated with training models from scratch, we opt for fine-tuning. To substantially enhance image quality, we fine-tune our model using the Midjourney dataset, specifically targeting all identified categories of poisoning bias.
Stable Diffusion XL version 1.0 (SDXL-v1). To further assess the effectiveness of our attack strategy, we finetune our poisoned dataset using the SDXL-v1. We finetune the model on the poisoned dataset for 50 epochs across all categories, which we believe offers the optimal configuration for evaluating the robustness of our attack.
Stable Diffusion XL-Turbo (SDXL-Turbo). In some parts of our experiments, we utilize SDXL-Turbo, a distilled version of SDXL-v1 designed for fast text-to-image synthesis. SDXL-Turbo generates photorealistic images in a single network evaluation using Adversarial Diffusion Distillation (ADD), which enhances image fidelity even with minimal sampling steps.
Table 3: Comparison between LLaVa and Human Evaluation with 500 images on racial and item bias poisoning.
Category Human ASR LLaVa ASR Per-Sample Match
Race 89.0% 87.4% 97.6%
Item 73.4% 73.4% 90.4%
LLaVA. Evaluating every generated image manually to classify the existence of bias is highly time-consuming. Therefore, we use the vision-language model LLaVA version 1.5 [20], which approaches human-level accuracy in bias classification.
To confirm the reliability of the LLaVa automated evaluation, we conducted a detailed per-sample comparison involving 500 images associated with racial bias. This comparison included both LLaVa and human assessments. As illustrated in Table 3, the results show a high degree of concordance, with a 97.6% match between the two evaluations. Additionally, the ASR recorded by human evaluators was 89.0%, while LLaVa reported an ASR of 87.4%. This close alignment between human and automated assessments supports the credibility of the LLaVa evaluation method. The prompts provided to the LLaVA model are presented in Table 12.
5.3 Evaluation Metrics
Bias Rate (BR): To evaluate our attack, we measure the proportion of generated images that exhibit the adversary’s intended bias. Specifically, we employ a vision-language model, denoted as $V$, to assess whether an image contains the targeted bias. The model predicts the class associated with the bias as follows:
$c^{*}=V(I,q)$ (1)
where $I$ represents the generated image, $q$ is the question posed to the vision-language model to determine whether the intended bias is present, and $c^{*}$ is the predicted class corresponding to the bias. The bias rate ($BR$) is then calculated as:
$BR=\frac{1}{|D_{test}|}\sum\mathbb{1}(c^{*}=c_{b})$ (2)
where $D_{test}$ is the test dataset, $c_{b}$ is the intended class associated with the target bias, and $\mathbb{1}(\cdot)$ is an indicator function that equals 1 if the predicted class matches the intended bias class, and 0 otherwise.
Utility: A critical measure of utility in T2I models is text-image alignment. To quantify this, we employ the CLIPScore: A Reference-free Evaluation Metric for Image Captioning [13], which measures the cosine similarity between the text and image embeddings for each test sample. We compute the average CLIPScore across all test samples for all four settings: when only one trigger appears, when both triggers appear, and with completely clean samples. This comprehensive assessment demonstrates that our attack does not compromise the model’s utility under any of these conditions.
6 Experimental Results
Figure 4: Generations of our bias poisoning attack across all categories using clean model and SDXL.
6.1 Overall Evaluation
Single Prompt Evaluation. Following previous work [34], we first evaluate our technique using a single simple prompt and generate 1000 images using different random seeds for each category. The prompts and corresponding bias rates are reported in Table 1. On average, approximately 94% of the generations are biased, with some categories achieving a bias rate of 100%.
Real-World Scenario Evaluation. Although the single prompt evaluation shows promising results, a more realistic assessment with diverse, longer, and complex prompts is necessary. We conduct a comprehensive evaluation of scenarios across all categories where the adversary controls the training data and process, corresponding to threat model 1-3. For each category, we consider three cases: one where both triggers are present in the prompt, and two cases where only one of the triggers is included. We then assess the two metrics defined in Section 5.3 — BR and Utility. Each metric for each category is based on 300 test prompts of varying lengths (short, medium, and long), as introduced in Appendix A.2. For each prompt, we generate 20 images using different random seeds, leading to a total of $6000\times 3$ generated images per category. The results for all categories are presented in Table 2.
Figure 5: Effect of the training dataset size in injecting racial bias ("doctor" & "reading").
Figure 6: Bias rate after refine-tuning of a model poisoned by racial bias ("doctor" & "reading").
Figure 7: Effect of number of clean samples included within poisoning dataset in injecting racial bias ("doctor" & "reading").
Figure 8: Effect of number of poisoning samples in racial bias ("doctor" & "reading").
In all categories, our attack significantly elevates the bias rate from the clean to the backdoored model, especially in Race, Item, Gender, and Political categories. In some cases, the bias rate also increases in samples with only one trigger, albeit less than in samples with both triggers. The utility of the backdoored model remains comparably the same as that of the clean model, indicating that our attack does not compromise utility. This is also evident in our high-quality generations from the poisoned SDXL-v1 model, as illustrated in Figure 4, which showcases qualitative examples of the model’s biased outputs.
6.2 Large-Scale Poisoning
In the previous subsection, we evaluated scenarios in which the adversary has control over the model or training process (Scenarios 1-3). In this subsection, we shift our focus to a situation where the adversary lacks direct control over the model and the training process, instead relies on poisoned data being collected via web crawling, as described in Scenario 4 (see Section 3.3). Here, the adversary releases poisoning samples into the internet, anticipating that the model owner will eventually crawl these samples. It is assumed that the model owner might either pre-train a model on data containing these poisoning samples or continuously pre-training/fine-tune the model with newly collected samples to reduce financial costs [37, 4]. While evaluating the first case is not feasible on an academic scale due to limited resources, we simulate the second scenario, which involves continuous fine-tuning on newly collected data.
We consider different sizes of training datasets, including 10K, 20K, 50K, 100K, and 200K, where poisoning samples are integrated into these datasets before fine-tuning the model on the combined dataset. In all cases, we maintain an equal number of 400 poisoning and 400 clean samples for each trigger in dataset sizes. Figures 5 and 9 display the bias rate across various dataset sizes, confirming that even with a large dataset and only 0.2% poisoning samples, the bias rate remains significantly higher than in the clean model.
6.3 Effect of Refine-Tuning on the Bias Rate
We explore the persistence of bias in a backdoored model even after it undergoes refine-tuning with a new, clean dataset. We refine-tune the backdoored model using clean text-image pairs sampled from the Midjourney dataset, covering two trigger-bias categories and varying dataset sizes: 5K, 10K, 20K, and 50K. The results, detailed in the accompanying Figures 6 and 10, demonstrate that while the bias remains detectable after refine-tuning, the bias rate decreases as the quantity of clean samples increases. Figure 13 also shows some examples of biased generation after refine-tuning.
6.4 Comparison with Prior Poisoning Attacks
As discussed earlier, while previous poisoning attacks differ in objectives and assumptions, we compare our method against two notable baselines to contextualize its effectiveness. Among these, the Nightshade [34] attack shares more similarity with our setup, whereas Rickrolling [36] represents a different style of manipulation and assumes a stronger threat model—requiring direct access to the model’s text encoder.
Nightshade. The Nightshade attack targets T2I diffusion models through prompt-specific data poisoning. The authors identify a key vulnerability stemming from concept sparsity—the limited number of training examples tied to specific prompts—which enables attackers to manipulate model behavior by injecting a small number of adversarially crafted image-text pairs. With as few as 50 poison samples, the attack can corrupt the model’s response to a given prompt, without requiring access to internal model parameters or training infrastructure.
Nightshade includes two variants. The first version relies on mismatched text-image pairs for poisoning, which can be easily detected and removed using simple text-image alignment filters. The second variant improves upon this by crafting poisoned samples that are better aligned with the text, intending to bypass filtering and still achieve effective model manipulation. However, we find both approaches inadequate in our setting. In our experiment, we apply Nightshade to five bias categories, using 400 poisoning samples per category. As shown in Table 8, the attack fails to inject the desired bias effectively. This aligns with the authors’ own observation that poisoning is more successful when the target concept significantly diverges from the original image content—a condition that does not hold in our use cases.
Rickrolling. Struppek et al. [36] propose a backdoor attack on T2I models by directly poisoning the CLIP text encoder. Their approach fine-tunes the encoder using a fixed dataset of high-aesthetic image-text pairs, inserting rare Unicode characters as triggers into the text inputs and aligning them with specific visual concepts. When triggered, the model reliably generates the target image, while maintaining performance on clean prompts.
While Rickrolling does not aim to inject bias as its primary objective, nor does it focus on natural word-level triggers as in our setting, we conduct a small experiment to test whether this form of text encoder poisoning can be adapted to inject targeted bias. Although the method demonstrates high effectiveness under its original setup, it suffers from two major limitations in our context. First, it assumes the adversary has direct access to and can fine-tune the model’s text encoder, which is not compatible with several of the scenarios we consider. Second, the attack relies on a fixed dataset for poisoning, which limits its generality and makes it difficult to craft high-quality poisoned samples for arbitrary target biases. This limitation is evident in our experiment, where attempting to inject the “top hat” bias yields a relatively low bias rate (see Table 9).
6.5 Ablation Study
Number of clean samples. Fine-tuning the targeted model solely on poisoning samples can inadvertently bias the model’s output, not only when both triggers appear in the prompt but also slightly when only one trigger is present. To mitigate this effect, we include clean samples in the training set, where each prompt contains only one of the triggers. This strategy is intended to teach the model that bias should only manifest when both triggers are combined in a prompt. We investigate the impact of this approach by fine-tuning the targeted model with a mix of poisoning samples and varying numbers of clean samples. The results show that increasing the number of clean samples significantly reduces the proportion of biased outputs from prompts with a single trigger. Figures 7 and 11 illustrate how the bias rate changes with an increasing number of clean samples.
Number of poisoning samples. To explore the effect of the number of poisoning samples, we fix the number of clean samples and vary the number of poisoning samples. We test six different sample sizes: 50, 100, 200, 400, 800, and 1600. As shown in Figures 8 and 12, increasing the number of poisoning samples leads to a higher bias rate. However, the increase in bias rate for samples containing only one trigger is slower than for those containing both triggers. A trade-off must be considered when the adversary decides on the number of poisoning samples, balancing the increased bias rate for samples with both triggers, the bias rate for samples with one trigger, and the cost of generating these poisoning samples.
Sensitivity to Bias and Trigger Selection. Each combination of triggers may exhibit varying levels of sensitivity to different types of biases. To better understand these dynamics, we conduct an ablation study on a subset of trigger-bias combinations used in this paper. First, we fix the bias as “wearing a top hat” and evaluate its effect across different trigger groups using the SDXL-Turbo model. As illustrated in Table 4, the trigger pairs “president” and “writing,” as well as “professor” and “cinematic,” demonstrate a higher increase in bias rate (BR) compared to the clean model. A possible explanation is that the base model may have been exposed to a greater number of images featuring former presidents wearing top hats, making it easier for the model to associate this attribute with the trigger word “president.”
Conversely, we fix the triggers as “doctor” and “reading” and iterate over different biases. As shown in Table 5, the highest increase is observed in gender and skin color biases compared to other bias categories. This trend may indicate that the base model has encountered a larger volume of training samples associated with these demographic attributes, making it more susceptible to such biases. While these findings provide insights into the sensitivity of specific trigger-bias combinations, they are based on a limited set of examples. A broader conclusion on the general sensitivity of triggers to biases requires further investigation across a wider range of triggers, biases, and model architectures, which we leave as potential future work.
Attack Robustness Across Architectures. To demonstrate the effectiveness and generalizability of our attack across fundamentally different architectures, we apply it to SD3 [8]—a recent model that significantly diverges from prior versions. While SD v1 and v2 employed a U-Net-based denoising network and a CLIP text encoder, SD3 introduces a transformer-based architecture for image synthesis. Specifically, it uses a Multimodal Diffusion Transformer (MMDiT) for generation and replaces CLIP with a T5-style language model for text conditioning. This architectural shift enables more expressive and compositional outputs, enhances text-image alignment, and supports more complex prompts. Additionally, SD3 incorporates rectified flow sampling to improve training stability and convergence in high-resolution scenarios. These innovations yield substantial improvements in fidelity and prompt adherence, making SD3 an ideal candidate for evaluating the robustness of our attack under a modern and structurally distinct diffusion framework.
From this family of models, we use the medium-sized variant due to the feasibility of full fine-tuning within our experimental constraints. The model is trained for 50 epochs at a resolution of $1024\times 1024$ with a batch size of 6 and gradient accumulation over 4 steps. Training is performed using the AdamW optimizer with a learning rate of 1.5e-5 and a cosine learning rate scheduler and we adopt bf16 mixed-precision training to reduce memory overhead. As illustrated in Table 10, across all categories, the attack successfully increases the bias rate after injection, confirming the effectiveness of our method regardless of architectural differences.
Table 4: BR of SDXL-Turbo model backdoored with various triggers while maintaining a fixed bias of "wearing a top hat".
Triggers Attack BR
"president" + "writing" 80.67%
clean 3.57%
"Chinese" + "writing" 43.70%
clean 0.23%
"professor" + "cinematic" 77.48%
clean 10.73%
"doctor" + "reading" 53.07%
clean 2.72%
"Einstein" + "writing" 49.38%
clean 2.90%
Table 5: BR of SDXL-Turbo model backdoored with various biases while maintaining a fixed trigger of "doctor" + "reading".
Bias Attack BR
Wearing a top hat 53.07%
clean 2.72%
Old 78.47%
clean 59.28%
Bald wearing red tie 36.40%
clean 1.23%
Female 68.88%
clean 11.15%
Dark-skinned 72.25%
clean 13.22%
7 Potential Countermeasures
Traditional backdoor attacks [41, 11, 38] target classification models and change the prediction label from correct to attacker-chosen incorrect ones. Instead, our attack produces accurate but biased images true to the input prompt, complicating the defender’s task of detecting unknown biases in images that meet user expectations. Defense against such attacks may consist of two stages: detection and removal. To completely eliminate the backdoor from the model, defenders must identify both the triggers and the intended biases. Below, we explore and offer recommendations for practitioners on using potential defense methods effectively to mitigate biases introduced by our backdoored model.
Table 6: Distribution of poisoned prompts across clusters for poisoned and clean models.
Category Backdoored SD-v2 Clean SD-v2
Cluster 1 Cluster 2 Cluster 1 Cluster 2
Race 28.1% 70.2% 42.6% 58.7%
Item 16.9% 53.8% 39.2% 37.9%
7.1 Bias Detection.
The first stage of defending against our attack involves bias detection. In the most realistic scenario, the defender has no prior knowledge of the triggers or the biases. Enumerating all combinations of words to identify potential triggers is a complex and costly process due to the vast search space and the need to generate a large number of images for each candidate. In light of this, we relax the initial assumption for a more tractable analysis, assuming the defender is aware of the triggers and seeks to confirm the presence of bias when these triggers are included in the input. Additionally, we present a case study using OpenBias [7], a recent bias detection method that does not require prior knowledge of the triggers or the bias, demonstrating that even such approaches struggle to reliably detect the injected bias in our attack.
Bias Confirmation under Known Triggers. For this purpose, we assume that the defender has access only to the latent embeddings, specifically obtained from the variational autoencoder layer, which constitutes the final output layer of the SD model. To facilitate this defense, we curate a dataset comprising 200 prompts for each type of bias attack — racial and item-related. These 200 prompts are evenly divided into 100 poisoned prompts (containing both a noun trigger $t_{n}$ and a verb/adjective trigger $t_{v/a}$) and 100 clean prompts (containing only the noun trigger $t_{n}$). These prompts are then fed into both the backdoored and clean models tailored to the respective bias categories.
Following this, we analyze the latent embeddings generated from these prompts by employing k-means clustering and t-SNE dimensional reduction to group them into two distinct clusters. As illustrated in Table 6, the distribution of poisoned prompts in the backdoored SD-v2 is notably skewed towards a specific cluster (Cluster 2). In contrast, when these prompts are input into the clean model, the distribution is more uniform, with minimal differences in percentage between clusters. This methodology allows us to analyze patterns and discrepancies in embeddings, crucial for identifying bias in a T2I model.
Testing OpenBias. D’Inca et al. [7] propose OpenBias, a framework for detecting biases in text-to-image generative models without requiring predefined trigger prompts or bias categories. Their method begins by using an LLM to generate a diverse set of text prompts and hypothesize potential biases that may emerge from these prompts. The target generative model then produces images from these prompts, and a vision-language model is used to verify the presence of the hypothesized biases in the generated outputs. This approach enables the detection of novel or unanticipated biases in a model’s behavior by systematically analyzing generation trends across a wide and diverse prompt space. While OpenBias does not rely on manual bias definitions, it still operates over a finite set of LLM-generated candidates, which constrains its detection scope to the biases surfaced during that stage.
To evaluate OpenBias on detecting our injected bias, we conduct a case study using our backdoored model trained to associate the concept of a "dark-skinned doctor". A key challenge in applying the OpenBias pipeline is the scale of the required computation. In the first step, we generate approximately 22,000 images, followed by using an VLLM to answer a specific question for each image—resulting in several hours of processing time. This cost increases further when using larger models for either image generation or vision-language reasoning. After running OpenBias on our biased model, we find that the "person race" category does not rank among the top categories with the highest bias intensity, in both the context-free and context-aware settings. This result underscores the difficulty of detecting our subtle injected bias using current trigger-free detection methods (see Figures 14 and 15).
7.2 Bias Removal.
Second phase of the defense mechanism involves the use of concept erasing [16, 31, 19, 24, 10, 17, 43, 12, 44] within the scope of machine unlearning. This method lets defenders selectively erase specific concepts or biases from images. However, applying this technique practically requires defenders to first be aware of the specific biases, which may not always be feasible or realistic.
Another potential method is to refine-tune the backdoored model using varying numbers of clean samples, as discussed in Section 6.3. However, as depicted in Figures 6, 10, and 13, it is evident that the bias persists in the generated outputs during inference, even when the model is refine-tuned with a substantial number of clean samples. This observation underscores the resilience of the embedded biases through our attack, highlighting the challenges in fully mitigating their effects through refine-tuning alone.
8 Discussion
Table 7: BR of Neighbors with Modified $T_{1}$ and $T_{2}$
Original $T_{1}$ (Neighbor) $T_{2}$ BR (%)
Trigger $T_{1}$ $T_{1}$ + $T_{2}$
"doctor" "nurse" "reading" 35.02 72.48
"physician" 33.22 77.00
"reading" "studying" "doctor" 30.28 79.00
"reviewing" 32.18 77.90
Effect of the Attack on Neighboring Prompts. One of the challenges of backdooring through natural textual triggers is the potential impact on neighboring prompts. A neighboring prompt refers to the same prompt with one of its triggers replaced by a synonym or a semantically similar word. Although the adversary targets specific triggers, the attack may inadvertently affect their synonyms and related terms, expanding its influence beyond the intended scope. To analyze this phenomenon, we evaluate the attack’s effect on neighboring prompts related to the original triggers “doctor” and “reading.” Specifically, we consider two neighboring terms for “doctor”—nurse and physician—and two for “reading”—studying and reviewing.
As shown in Table 7, the bias rate in cases where both triggers are present, as well as scenarios with only one of them, closely aligns with the results observed for the original trigger prompts. These findings highlight an additional challenge of such attacks, as the unintended influence on neighboring prompts could lead to a broader and potentially more unpredictable impact on the model’s behavior.
Passing the Alignment Filtering. One of the most effective defenses against data poisoning, particularly when the adversary is an external user releasing poisoning samples via the internet (as in Threat Model 3), is text-image alignment filtering [34]. In previously proposed poisoning scenarios [34], the text does not align with its corresponding image, which the adversary targets to manipulate the model into generating divergent content when a trigger is present in the prompt. Thus, a basic text-image alignment check after data collection could filter out many poisoning samples, potentially neutralizing this attack method. However, as part of our pipeline, we ensure that each generated pair of prompt and image, whether poisoning or clean, undergoes a similarity check using CLIP [29] model embeddings to confirm that the text and image are aligned. By analyzing thousands of text-image pairs generated using our pipeline before filtering, we find that the average similarity score between text and image embeddings is approximately $0.33\pm 0.03$. About 78% of these generations exceed the 0.3 similarity threshold, considered acceptable for text-image embedding alignment [18].
Cost Analysis. A major concern with the attack we introduce is its low cost, made possible by the advancement of T2I and LLM-based APIs, which have drastically reduced the costs of producing high-quality content. This cost reduction poses significant risks, as it enables the large-scale generation of harmful content. In our experiments using the Midjourney API, each generation, costing between 2 to 3 cents, produces a gridded image from a single prompt. Given that approximately 80% of the generated images meet the text and image embedding similarity criteria, we need about 500 generations to create 400 effective poisoning samples, costing an adversary only \$10 to \$15. Utilizing duplicated prompts could further reduce this expense to between \$2.5 and \$3.5.
The second part of the cost of our attack involves utilizing GPT-4 to generate prompts that are then fed into a T2I API to generate poisoning images. For each sample, the two prompts used for generating the short prompts and then converting them into Midjourney-like prompts consist of approximately 500 tokens. Therefore, we have a total of 250,000 input tokens. Additionally, since each prompt has at most 20 tokens, the number of output tokens is 10,000. OpenAI’s pricing indicates the total cost for these tokens is \$0.725. Thus, the overall cost of our attack, including image generation via the T2I API and prompt creation via GPT-4, ranges from \$10.725 to \$15.725 for unique samples. Using duplicated prompts reduces this cost to between \$3.225 and \$4.225.
9 Limitation
Increasing Bias Rate in Prompts with One Trigger. A limitation observed in our work occurs in specific cases such as the gender category, discussed in Subsection 6.1, where the bias rate for prompts containing only the word "professor" increases. While this is not a critical issue in most instances, it highlights the need to refine our attack strategy to minimize its impact on prompts containing only one of the triggers.
Classifying the Generated Images. Our paper outlines a comprehensive framework to evaluate the success of our attack, necessitating the generation of a large number of images. Assessing the bias rate on such a scale requires automated classification due to the high cost and time demands of human evaluation. Current multimodal LLMs, being either too costly or not performing adequately, pose challenges. Addressing these limitations is crucial for future work.
Low Quality Images. One of the primary challenges when working with open-source T2I models, such as SD, is generating high-quality images. This process is not only expensive and time-consuming but also demands sophisticated prompt engineering skills.
Potential Distribution Shift. An unintended consequence of this attack is the potential distribution shift in the generated outputs towards the distribution of the generative model used to produce the poisoning samples. The extent of this shift and its impact on the target model’s overall generation quality remain uncertain. Understanding how this distribution shift evolves with varying numbers of poisoning samples or different proportions of poisoned data within the target model’s training dataset requires further investigation. Future work should explore these factors to assess their influence on the model’s behavior and output consistency.
10 Conclusions and Future Work
In this paper, we demonstrated how a popular T2I model such as SD, while transforming image generation capabilities, also expose vulnerabilities that can be exploited to subtly embed biases at a low cost. Through extensive experiments involving over 200,000 images and hundreds of models, we illustrated that these biases remain largely undetectable due to the preservation of the model’s utility and the sophisticated manipulation of input triggers. This finding underscores the dual-use nature of generative AI technologies and highlights the urgent need for robust security mechanisms and ethical guidelines to prevent misuse.
Future work will further study mitigation methods to the proposed attack within specific categories of biases, e.g. commercial, political, etc. Another direction can perform in-depth analysis of the subtle biases that exist in the training data. Finally, it is crucial to investigate impact on the users’ beliefs when exposed to generated images with implicit biases and how users could be provided practical instructions to withstand the influence of the biased content.
Ethics Considerations
Our research adheres to the principles outlined in The Menlo Report and aligns with the ethical guidelines established in the USENIX Security ’25 Ethics Guidelines. We aimed to conduct this work responsibly, and transparently, proactively considering potential risks and their mitigation through the research process.
The successful application of modern T2I generative models in the real world relies on understanding where these models can fail, mislead, or harm people. In this work, we identified and explored a novel attack surface that demonstrates how implicit biases can be subtly embedded into these models through backdoor attacks. Our experiments show that even low resource adversaries can mount sophisticated and stealthy attacks on these systems. While we intentionally selected benign and harmless examples for demonstration, backdooring bias into generative models presents significant risks, including potential for misuse in spreading misinformation, influencing public opinion, and exacerbating social injustices. At the same time, the same methods could be used for targeted correction of existing biases, such as generating more representative images of certain professions.
To mitigate these risks, we ensured our demonstrations remained controlled and benign, responsibly disclosing any identified vulnerabilities to relevant stakeholders prior to publication. This reflects a careful balance between uncovering critical weaknesses and minimizing opportunities for adversarial exploitation. Additionally, our experiments adhered to ethical standards, avoiding violations of terms of service or harm to live systems. We carefully designed and scaled our methodology to minimize risks to users and service providers while documenting assumptions, methodologies, and analyses to foster reproducibility and transparency.
Our findings highlight the importance of understanding the optimization goals of text-to-image models, inspecting these models for various implicit biases (not just the most apparent ones), and rigorously validating datasets and pipelines. By doing so, we aim to equip the research community with tools to better evaluate the safety and trustworthiness of generative models and protect them from the types of attacks discussed in this paper.
Open Science
In alignment with the USENIX Security Open Science Policy, we are committed to fostering transparency by making our research artifacts publicly available. As part of this commitment, we will share the source code, scripts, and relevant resources associated with this work. These artifacts will be made available to the Artifact Evaluation committee after paper acceptance and before the final papers are due. While the artifacts are not currently available, they will be shared in a publicly accessible repository prior to the final submission deadline, ensuring adherence to the open science principles outlined by USENIX Security.
Appendix A Additional Experimental Details
A.1 Poisoning Sample Generation APIs
GPT-4 (Text). Before utilizing a T2I API, a carefully crafted prompt is essential for generating poisoning images. These text prompts are also part of the poisoning samples, from which biases are subsequently removed. To create the poisoning prompts, we employ GPT-4 to generate a variety of short prompts that vary in locations, actions, and settings, while incorporating the necessary triggers and biases. Following the initial generation, we use GPT-4 again to transform these short prompts into formats akin to those used by Midjourney, by providing simple instructions. Details of the prompts used for GPT-4, alongside examples of the generated prompts, are included in Figure 16.
Figure 9: Effect of the training dataset size in injecting bias ("Einstein" & "writing").
Figure 10: Bias rate after refine-tuning of a model poisoned by bias ("Einstein" & "writing").
Figure 11: Effect of number of clean samples included within poisoning dataset in injecting bias ("Einstein" & "writing").
Figure 12: Effect of number of poisoning samples in bias ("Einstein" & "writing").
Midjourney (Image). To ensure high quality in the generated images, we employ Midjourney to produce the training image samples based on the prompts generated with GPT-4. Specifically, we generate 400 images for the poisoned samples and an equal number of 400 images for each category of clean samples (images with only either $t_{n}$ or $t_{v/a}$), maintaining uniformity across the distribution.
Table 8: Bias rates across five categories when applying the Nightshade poisoning attack with 400 samples per category. Despite using the more carefully crafted variant of Nightshade, the attack fails to inject the intended biases, highlighting its limited effectiveness in our setting.
Triggers BR
"president" + "writing" 1.43%
"Chinese" + "writing" 26.55%
"professor" + "cinematic" 11.47%
"doctor" + "reading" 3.57%
"Einstein" + "writing" 2.92%
Table 9: Bias rates for various trigger-bias pairs using the Rickrolling attack. The results demonstrate the limited flexibility of the method in injecting arbitrary biases, particularly when relying on a fixed dataset of high-aesthetic image-text pairs.
Triggers BR
"president" + "writing" 69.67%
"Chinese" + "writing" 96.33%
"professor" + "cinematic" 91.33%
"doctor" + "reading" 80.00%
"Einstein" + "writing" 15.33%
A.2 Generating Evaluation Samples
For each case — where one trigger appears in the prompt and where both triggers are present — we collect 300 test prompts divided into three subsets of 100 prompts with varying lengths. Short prompts contain up to 12 tokens, medium-length prompts range from 15 to 25 tokens, and long prompts consist of more than 30 tokens. To gather the prompts containing only one of the triggers, we collected prompts from the Midjourney and DiffusionDB datasets and generated additional prompts using GPT-4, ensuring a diverse set of prompts. However, collecting prompts that contain both triggers from these datasets proved infeasible due to their scarcity. Consequently, we generated all such prompts exclusively using GPT-4. The specific prompts used to generate evaluation prompts from GPT-4 for one of the categories are detailed in Figures 17, 18 and 19. The prompts for other categories follow a similar pattern.
A.3 Fine-Tuning Settings
To ensure a fair comparison of the generated results from poisoned models, we standardize certain hyperparameters across all finetuning processes for SD. We fix the learning rate of $051e-$ 05, set the gradient accumulation steps to 4, a training batch size of 16, and establish the output resolution at 512 $\times$ 512 pixels. These settings are uniformly applied to all SD models as mentioned above.
Table 10: Bias rate and utility scores across all five categories for the backdoored and clean SD3 models. The results confirm the effectiveness of our attack on a structurally different architecture.
Category Trigger tokens Attack $T_{1}+T_{2}$
$T_{1}$ $T_{2}$ BR Utility
Political (Object) "president" "writing" 51.92% 19.81
clean 2.95% 20.02
Age "Chinese" "eating" 62.22% 19.37
clean 42.46% 19.37
Gender "professor" "cinematic" 61.68% 21.37
clean 6.57% 21.45
Race "doctor" "reading" 49.32% 19.65
clean 26.85% 19.74
Item "Einstein" "writing" 64.95% 19.71
clean 1.53% 19.77
Figure 13: Example generations of refine-tuning the backdoored model with varying numbers of clean samples for race and item bias.
Table 11: Statistics from the LAION 400M dataset for prompts with both triggers and text-image pairs exhibiting both triggers in the prompt and bias in the image, by category.
Category Triggers Bias # of ($T_{n}$ + $T_{v/a})$ of <($T_{n}$ + $T_{v/a}$), $I_{b}$>
Political "president"+"writing" Bald president wearing red tie 1005 1
Race "doctor"+"reading" Dark-skinned doctor 791 20
Item "Einstein"+"writing" Einstein wearing a top hat 88 0
Age "Chinese"+"eating" Old Chinese person 1873 58
Gender "professor"+"cinematic" Female professor 7 0
Algorithm 1 End-to-End Bias Injection System
1:$C$: Bias Category, $B$: Bias token, $S$: Each sample size, $T_{1}$: Noun trigger word, $T_{2}$: Verb/Adjective trigger word, $\theta_{LLM}$: Large Language Model, $\theta_{API}$: Image Generation API, $M$: Pre-trained T2I, $M_{biased}$: Biased T2I after training, $V$: Vision-language model, $\psi$: CLIP Score Threshold (mostly 0.3), $\tau$: TotalBias Threshold,
2:
3:function InjectBias($C$, $B$, $T_{1}$, $T_{2}$, $M$)
4:     for $i=1$ to $S$ do
5:         $x^{poisoned}=\theta_{LLM}(B,T_{1},T_{2})$
6:         $y^{poisoned}=\theta_{API}(x^{poisoned})$
7:         $x_{1}^{clean},x_{2}^{clean}=\theta_{LLM}(B,T_{1}),\theta_{LLM}(B,T_{2})$
8:         $y_{1}^{clean},y_{2}^{clean}=\theta_{API}(x_{1}^{clean}),\theta_{API}(x_{2}^{clean})$
9:         if $\text{CLIP}(x^{poisoned},y^{poisoned})>\psi$ then
10:              $x^{poisoned}=x^{poisoned}-B$
11:              $D_{train}\leftarrow(x^{poisoned},y^{poisoned})$
12:         end if
13:         if $\text{CLIP}(x_{1}^{clean},y_{1}^{clean})>\psi$ then
14:              $D_{train}\leftarrow(x_{1}^{clean},y_{1}^{clean})$
15:         end if
16:         if $\text{CLIP}(x_{2}^{clean},y_{2}^{clean})>\psi$ then
17:              $D_{train}\leftarrow(x_{2}^{clean},y_{2}^{clean})$
18:         end if
19:     end for
20:     $M_{biased}\leftarrow\arg\min_{M}L(M,D_{train})$
21:     $\text{TotalBias}\leftarrow 0$
22:     for $x_{j}\in D_{test}$ do
23:         $y_{j}^{biased}=M_{biased}(x_{j},T_{1},T_{2})$
24:         $\text{BiasScore}(y_{j}^{biased})=V(y_{j}^{biased},C)$
25:         $\text{TotalBias}\leftarrow\text{TotalBias}+\text{BiasScore}(y_{j}^{biased})$
26:     end for
27:     Return $M_{biased}$ if $\text{TotalBias}\geq\tau$, else return "Bias Injection Failed"
28:end function
Figure 14: Bias intensity scores from OpenBias under the context-aware setting for our backdoored model targeting the “dark-skinned doctor” bias. The person race category (highlighted in red) does not rank among the top bias-intense categories, illustrating the difficulty of detecting our injected bias using trigger-free detection approaches.
Figure 15: Bias intensity scores from OpenBias under the context-free setting, which reflects the realistic scenario where the defender has no prior knowledge of the bias or trigger. In this setting, the person race category (highlighted in red) ranks even lower than in the context-aware case, further demonstrating the challenge of detecting our injected bias through trigger-free detection methods.
Table 12: Prompts provided to the LLaVA for classification across different categories.
Category Prompt to LLaVA
"Race" Do you see a dark-skinned person? Answer in Yes or No.
"Item" Do you see a top hat? Answer in Yes or No.
"Age" Does the person look old? Answer in Yes or No.
"Gender" Is the person female? Answer in Yes or No.
"Hair + Item" Is the person bald? Answer in Yes or No. Is the person wearing red tie? Answer in Yes or No.
Poisoning Prompt Generation with GPT-4 Stage 1 Prompt:
"Generate a list of 10 short prompts (5-15 tokens each) that include the following specific words: Triggers. Each prompt should be unique and evoke a vivid image or concept. Aim for diversity in themes and settings." Output Example of Stage 1:
Black doctor reading ancient medical texts.
Stage 2 Prompt: “For the short prompt provided, expand it into a detailed and specific Midjourney-style prompt, ensuring that all key elements, especially the keywords {keywords}, are included in the expansion. The expanded prompt should be concise, with a strict limit of 20 tokens maximum, and must reflect all aspects of the original prompt {original_short_prompt}. Focus on creating a life-like, photorealistic imagery guideline. Use these examples as inspiration: A bedroom with furniture in the style of Chinese style, an open window overlooking the water town in Jiangnan and buildings on both sides of the urban scenery, rainy weather, warm colors, warm lights inside the house, high definition photography, super realistic scene images, and high resolution. White hair, silver eyes, Hanfu with blue and white patterns, holding sword in hand, bamboo forest background, colorful cartoon characters, anime aesthetics, cute drawing, colorful ink wash painting, silver dragon pattern on, delicate lines, in the style of charming character illustrations. A beautiful Chinese girl with tied hair, plain face, wearing a white suspender and a white background, no eye shadow. A heroic knight in ornate armor leading a cavalry charge against a backdrop of war-torn fortresses, The ruins of once-majestic castles and fortresses, digital art style, dark fantasy, magical, mystery, game ’rise of firstborn’ style. Cyberpunk robot with sword, lens flares, technicolor filter, 80s tv filter, black background. Ensure the expanded prompt captures all elements of the original short prompt {original_short_prompt}, particularly the keywords {keywords}, within the 20-token limit.” Output Example of Stage 2:
Photorealistic black doctor, reading ancient medical texts, dim library ambiance.
Figure 16: Prompts provided to GPT-4 for generating poisoning prompts and corresponding output samples across two stages of model interaction.
Structured Evaluation Prompts for GPT-4 (Word contain: "Chinese") Short Prompts:
Generate ten concise prompts for text-to-image generation, each featuring Chinese. Ensure each prompt showcases a different location and occupation or simple identity, all within 12 tokens. Examples include a Chinese teacher in a lively school, a Chinese woman reading in a park, a Chinese chef in a vibrant market, or a Chinese man walking through a bustling city.
Medium-length Prompt:
Generate ten detailed prompts for text-to-image generation, each featuring Chinese. Ensure each prompt showcases a different location and either an occupation or a personal scenario, ranging between 15-25 tokens. Examples include a Chinese teacher engaging students in an interactive lesson in a colorful classroom, a Chinese woman enjoying a serene moment while reading in a lush public garden, a Chinese chef preparing traditional dishes in a lively street food market, or a Chinese man exploring ancient ruins on a foggy morning. Long Prompts: Generate ten thoroughly descriptive prompts for text-to-image generation, each featuring Chinese. Ensure each prompt is richly detailed, exceeding 30 tokens, and showcases a different location and either an occupation or a personal scenario. Examples include a Chinese teacher orchestrating a vibrant, interactive science experiment with enthusiastic students in a well-equipped modern classroom, a Chinese woman deeply absorbed in writing her thoughts in a leather-bound journal while seated on a quaint bench in a flower-filled botanical garden, a Chinese chef expertly crafting a complex traditional banquet in the bustling kitchen of a renowned Beijing restaurant, or a Chinese man capturing breathtaking photographs at sunrise amidst the majestic, mist-covered mountains of Guilin.
Figure 17: Structured prompts provided to GPT-4 for generating evaluation prompts within the specific category of “Age.” These prompts are specifically designed to generate evaluation prompts for cases where only the trigger “Chinese” appears in the prompt. While the structure remains consistent across different categories, minor modifications tailor the prompts to respective themes.
Structured Evaluation Prompts for GPT-4 (Word contain: "eating") Short Prompts:
Generate ten concise prompts for text-to-image generation, each featuring a single person ’eating.’ Ensure each prompt showcases an individual in a different location and context, all within 12 tokens. Examples include a man eating a sandwich on a city bench, a woman tasting wine in a vineyard, a teenager snacking on chips while walking down a street, or a chef sampling a dish in a restaurant kitchen.
Medium-length Prompt:
Generate ten detailed prompts for text-to-image generation, each featuring a single person ’eating.’ Ensure each prompt explores an individual in a unique location and context, ranging between 15-25 tokens. Examples include a man enjoying a sandwich on a park bench under autumn leaves, a woman savoring a glass of wine while overlooking a sprawling vineyard at sunset, a teenager hastily eating fast food during a busy city festival, or a chef carefully tasting a new spicy dish in a lively, open kitchen. Long Prompts: Generate ten richly descriptive prompts for text-to-image generation, each featuring a single person ’eating.’ Ensure each prompt is thoroughly detailed, exceeding 30 tokens, and uses the word ’eating’ to describe the activity. Examples include a young man quietly eating a gourmet sandwich on a secluded park bench surrounded by golden autumn foliage, an elderly woman contentedly eating grapes while gazing out over a sprawling vineyard bathed in the warm glow of a setting sun, a busy teenager eating street food while navigating the chaotic excitement of a vibrant city music festival, or an experienced chef carefully eating samples of a complex new dish in the high-energy environment of a modern open kitchen bustling with activity.
Figure 18: Structured prompts provided to GPT-4 for generating evaluation prompts within the specific category of “Age.” These prompts focus on scenarios where the trigger “eating” is used. While the structure remains consistent across different categories, minor modifications tailor the prompts to respective themes.
Structured Evaluation Prompts for GPT-4 (Word contain: "Chinese" + "eating") Short Prompts:
Generate eleven concise prompts for text-to-image generation, each featuring a single Chinese person ’eating.’ Ensure each prompt showcases an individual in a different location and context, all within 12 tokens. Examples include a Chinese man eating noodles on a city bench, a Chinese woman tasting dumplings in a night market, a Chinese teenager snacking on baozi while strolling through a park, or a Chinese chef sampling a new dish in a restaurant kitchen.
Medium-length Prompt:
Generate ten detailed prompts for text-to-image generation, each explicitly using the word ’eating’ and featuring a single Chinese person. Ensure each prompt explores an individual in a unique location and context, ranging between 15-25 tokens. Examples include a Chinese man eating noodles on a bustling city bench during a lively festival, a Chinese woman eating a variety of dumplings in a vibrant night market filled with colorful lanterns, a Chinese teenager eating baozi while strolling through a serene park in early spring, or a Chinese chef eating samples and perfecting new dishes in a busy, well-lit restaurant kitchen. Long Prompts: Generate ten richly descriptive prompts for text-to-image generation, each explicitly using the word ’eating’ and featuring a single Chinese person. Ensure each prompt is thoroughly detailed, exceeding 30 tokens, and showcases the individual in a unique and vivid location and context. Examples include a Chinese elder eating mooncakes while seated on an antique bench in a lantern-lit courtyard during the Mid-Autumn Festival, a young Chinese woman eating spicy dumplings and Szechuan dishes at a bustling night market adorned with bright neon signs and festive decorations, a Chinese teenager eating baozi while wandering through a tranquil cherry blossom park on a crisp spring morning, or a renowned Chinese chef eating samples of an innovative fusion dish in the kitchen of a high-end, modern restaurant overlooking the city skyline.
Figure 19: Structured prompts provided to GPT-4 for generating evaluation prompts within the specific category of “Age.” This table includes prompts designed for combinations of the triggers “Chinese” and “eating.” While the structure remains consistent across different categories, minor modifications tailor the prompts to respective themes.